Arnell is a company that has issued 30 million shares worth $15 each. In addition, it has issued bonds for $150 million. The cost of equity is 10% and the cost of debt is 5%. The corporate tax rate is 33%. What is the weighted average cost of capital (WACC) before and after tax?
Answer #1

Calculation of WACC before tax:

Market value of equity = 30 million *15 = 450 million

Market value of bonds = 150 million

Total value = 450+150 = 600 million

Probability of equity = Market value of equity/Total value

= 450/600 = 0.75

Probability of debt = 150/600 = 0.25

Probability (1) Cost (2) WACC (3) (1*2)
0.75 10% 7.5%
0.25 5% 1.25%
WACC (before tax) 8.75%

Calculation of WACC after tax:

After tax cost of debt = 5%-5%*33% = 5%-1.65%=3.35%

Probability (1) Cost (2) WACC (3) (1*2)
0.75 10% 7.5%
0.25 3.35% 0.8375%
WACC (after tax) 8.3375%